# Quantile Reward Policy Optimization: Alignment with Pointwise Regression and Exact Partition Functions

### v1

@@ 6 Conclusion and discussion
In this work, we have presented Quantile Reward Policy Optimization (QRPO), a novel RL fine-tuning algorithm that uses quantile rewards to derive a tractable expression of the optimal solution to the KL-regularized RL objective with its partition function, and fits it with a simple regression on individual samples; therefore, directly using the pointwise absolute reward of samples instead of relying on preferences. We have shown that QRPO provides strong empirical results on chat and coding tasks, scales with the pre-computation budget, and is less prone to length bias. We believe that QRPO unlocks a new avenue for policy fitting methods, which can combine any data distribution during training, benefit from offline scaling, and now use the signal from absolute pointwise rewards.
@@
We acknowledge the limitations of empirical research and that variations in results may stem from differing experimental conditions. Still, we strived for a fair and comprehensive evaluation using sufficient datasets, models, and hyperparameters, with equal budgets and statistical rigor within our compute limits. QRPO provides a solution to obtain a tractable expression of the partition function using quantile rewards, which turned out to be very effective in our experiments; however, the quantile reward reshapes the reward distribution, so for some tasks, the new shape may not be optimal. Nevertheless, we have proposed QRPO as a framework and shown that transformations can be applied on top of the quantile to reshape it into the desired form while still maintaining a tractable partition function. We therefore encourage future work to study the transformations that can be applied on top of the quantile reward. We provide guidelines in Appendix B. Perhaps another limitation of the quantile reward is that it is based on the reference policy, and when the model improves drastically beyond the reference policy, the quantiles become less informative. Still, we believe this can be overcome by performing iterative training and updating the reference policy, as is done in many state-of-the-art open-source post-training pipelines. Finally, QRPO and policy fitting methods in general can leverage any distribution, including online data, but we only focused on offline data to compare with other policy fitting methods. An open question is the impact of online data, potentially in combination with offline and off-policy data, and a comparison to policy improvement methods such as GRPO, which can only use online data.
@@ Appendix A QRPO loss and gradient interpretation
or with an additional transformation of the quantile reward ~$\tilde{\mathcal{R}}=f(\mathcal{R}_{q})$ and its corresponding $\tilde{Z}$ as described in Equation 13 and Appendix B.
@@
$-\nabla_{\theta}\mathcal{L}_{QRPO}=2\beta\mathop{\mathbb{E}}_{x,y}\left[\left(\bigg{[}\underbrace{\mathbf{\mathcal{R}}_{q}(x,y)-\beta\log{Z_{q}}}_{\textit{calibrated target}}\bigg{]}-\beta\log{\frac{\pi_{\theta}(y\mid x)}{\pi_{ref}(y\mid x)}}\right)\nabla_{\theta}\log\pi_{\theta}(y\mid x)\right].$ (14)
@@
Figure 6: $\log{\frac{\pi_{\theta}(y\mid x)}{\pi_{ref}(y\mid x)}}$ dynamics for chosen and rejected responses when training with QRPO on a preference dataset in the offline regime with $\beta=0.1$. We plot the average completion $\log$-ratio over all the prompts in the evaluation set. $\log$-ratios start at 00 when the trained policy is initialized $\pi_{\theta}(y\mid x)=\pi_{ref}(y\mid x)$, and during training become positive if $\pi_{\theta}(y\mid x)>\pi_{ref}(y\mid x)$ and negative if $\pi_{\theta}(y\mid x)<\pi_{ref}(y\mid x)$. As desired, we observe that they become positive for chosen completions and negative for rejected completion. Indeed, with $\beta=0.1$ the quantile threshold for samples to increase their probabilities is at $\mathcal{R}_{q}\approx 0.77$ and in this dataset chosen samples have an average $\mathcal{R}_{q}$ of approximately 0.91, while rejected samples have an average of around 0.45. The initial decrease of the chosen $\log$-ratios can be attributed to generalization caused by the gradient being larger for rejected completions, until a clearer separation between them is learned by the model.
@@
Interpretation of the gradient in the online case
@@ Appendix B Guidelines for using QRPO with quantile transformation functions
[…]
@@
~$\mathbf{\tilde{\mathcal{R}}}(x,y)=f\left(\mathbf{\mathcal{R}}_{q}(x,y)\right).$
@@
The optimal solution for the KL-regularized RL fine-tuning objective (Eq. 2) depends on the shape of the reward distribution; therefore, the primary function of a quantile transformation $f$ is to modify the transformed reward distribution to have some useful properties. One can easily obtain any desired distribution by using its inverse CDF as a quantile transformation function (since for any distribution $Y$, applying its inverse CDF to a uniform sample, $x\sim\texttt{Uniform}[0,1]$, results in $\texttt{CDF}_{Y}^{-1}(x)\sim Y$). For example, one might wish the reward distribution to be Gaussian with mean $\mu$ and variance $\sigma^{2}$. However, after transforming it into a quantile reward $\mathcal{R}_{q}$, the resulting reward is uniformly distributed. In this case, the following transformation can be applied to recover the desired distribution: ~$\tilde{\mathcal{R}}=f(\mathcal{R}_{q})=\sigma\cdot\texttt{CDF}^{-1}_{\mathcal{N}(0,1)}(\mathcal{R}_{q})+\mu$.
@@
One possible approach to address the finiteness condition is to get insights from the distribution of transformed rewards $\mathbf{\tilde{\mathcal{R}}}$. The partition function $\tilde{Z}(x)$ is simply a moment generating function (MGF) for the distribution of transformed rewards under the reference model: $\tilde{Z}(x)=\mathbb{E}_{\tilde{r}}\left[e^{t\tilde{r}}\right]=M_{\tilde{r}}(t)$, where $t=\frac{1}{\beta}$, $\tilde{r}=\mathbf{\tilde{\mathcal{R}}}(x,y)$ for $y\sim\pi_{ref}(\cdot\mid x)$ (see Section C for the derivation). Thus, to check the finiteness of $\tilde{Z}(x)$, one can derive the distribution of the transformed rewards for completions sampled from the reference model, and then check the value of the corresponding MGF for $t=\frac{1}{\beta}$.
@@
There is a sufficient condition for the finiteness of $\tilde{Z}$: if the transformation function $f$ is upper-bounded on the interval $[0,1]$, then the corresponding partition function is guaranteed to be finite (proof in Appendix F). However, it is very important to note that this condition is not necessary, i.e., there exist some unbounded functions (we consider the upper bound) such that the corresponding $\tilde{Z}$ is finite. An important example of such an unbounded function is the inverse CDF for the normal distribution (the transformation we considered above to obtain a Gaussian distribution of transformed rewards).
@@
After applying a quantile transformation to the initial reward values, the distribution of the resulting quantile reward $\mathcal{R}_{q}$ is known; therefore, the corresponding partition function $Z_{q}$ can be calculated analytically. Moreover, starting from a quantile reward $\mathcal{R}_{q}$, one can apply any additional transformation $f(\mathcal{R}_{q})$ to it while still preserving the ability to compute the corresponding partition function: the partition function corresponding to a specific transformation function $f$ can be computed as a simple integral (derivation in Appendix E):
[…]
@@
We can make a more rigorous observation regarding the correct scaling of $\beta$ in the case where the quantile transformation function $f$ is upper-bounded on the interval $[0,1]$ (see Appendix J for the intuition). The key factor determining the relative scale of the transformed reward is the left derivative of the quantile transformation function $f$ at $x=1$. Therefore, when comparing two upper-bounded quantile transformation functions, the ratio between their respective $\beta$ values should be approximately equal to the ratio of their left derivatives at $x=1$. If both left first derivatives at $x=1$ are equal to zero, move to higher orders until you find the first non-zero left derivative common to both functions. Let $k$ be that order. Then the appropriate $\beta$-scaling factor is the ratio of their $k$-th left derivatives. If the first non-zero left derivative appears at different orders for the two functions, no single constant can scale $\beta$ consistently.
Appendix C The partition function is a moment generating function at $t=1/{\beta}$

### v2

@@ 6 Conclusion and discussion
We have presented Quantile Reward Policy Optimization (QRPO), a novel RL fine-tuning algorithm that uses quantile rewards to derive a tractable expression of the optimal solution to the KL-regularized RL objective with its partition function, and fits it with a simple regression on individual samples; therefore, directly using the pointwise absolute reward of samples instead of relying on preferences. We have shown that QRPO provides strong empirical results on chat and coding tasks, scales with the pre-computation budget, and is less prone to length bias. We believe that QRPO unlocks a new avenue for policy fitting methods, which can combine any data distribution during training, benefit from offline scaling, and now use the signal from absolute pointwise rewards.
@@
We have proposed QRPO as a framework and shown that transformations can be applied on top of the quantile to reshape it into the desired form while still maintaining a tractable partition function. However, we have not explored these transformations empirically. We encourage future work to study them and provide further understanding in Appendix B. A notable limitation of the quantile reward is that it is based on the reference policy, and when the model improves drastically beyond the reference policy, the quantiles become less informative. Still, this can be overcome by performing iterative training and updating the reference policy, as is done in many state-of-the-art open-source post-training pipelines. Finally, QRPO and policy fitting methods in general can leverage any distribution, including online data, but we only focused on offline data to compare with other policy fitting methods. An open question is the impact of online data, potentially in combination with offline and off-policy data, and a comparison to policy improvement methods such as GRPO, which rely critically on online data.
@@ Appendix A QRPO loss and gradient interpretation
or with an additional transformation of the quantile reward $\tilde{\mathcal{R}}=f(\mathcal{R}_{q})$ and its corresponding $\tilde{Z}$ as described in Equation 13 and Appendix B.
@@
[…]
@@
Figure 6: $\log{\frac{\pi_{\theta}(y\mid x)}{\pi_{ref}(y\mid x)}}$ dynamics for chosen and rejected responses when training with QRPO on a preference dataset in the offline regime with $\beta=0.1$. We plot the average completion $\log$-ratio over all the prompts in the evaluation set. $\log$-ratios start at 0 when the trained policy is initialized $\pi_{\theta}(y\mid x)=\pi_{ref}(y\mid x)$, and during training become positive if $\pi_{\theta}(y\mid x)>\pi_{ref}(y\mid x)$ and negative if $\pi_{\theta}(y\mid x)<\pi_{ref}(y\mid x)$. As desired, we observe that they become positive for chosen completions and negative for rejected completion. Indeed, with $\beta=0.1$ the quantile threshold for samples to increase their probabilities is at $\mathcal{R}_{q}\approx 0.77$ and in this dataset chosen samples have an average $\mathcal{R}_{q}$ of approximately 0.91, while rejected samples have an average of around 0.45. The initial decrease of the chosen $\log$-ratios can be attributed to generalization caused by the gradient being larger for rejected completions, until a clearer separation between them is learned by the model.
@@
Interpretation of the gradient in the on-policy case
@@
On-policy training further improves QRPO performance
Figure 7: Performance of QRPO under two training regimes: using only offline completions and using a mix of offline and online completions.
It is well established that on-policy training improves the performance of policy optimization methods (lanchantin2025bridging; guo2024direct), as it allows for better control over output probability mass drift. Although on-policy training is beyond the main scope of this work, we nevertheless verify that QRPO also benefits from this regime. Specifically, we conduct ablations where online data is incorporated alongside offline data during QRPO training on the Magpie-Air dataset. This results in a 26% greater reward improvement compared to training with offline data alone (see Figure 7), highlighting the strong potential of QRPO in on-policy settings.
Dynamic-reference KL for stable and flexible QRPO
QRPO inherently optimizes a KL-regularized reinforcement learning objective (Eq. 1). While there is ongoing debate about the necessity of explicit KL penalties, particularly in the context of verifiable rewards (yu2025dapo), recent work presents a more nuanced perspective: liu2025prorl demonstrate that combining KL regularization with periodic updates of the reference policy to recent checkpoints both stabilizes training and prevents stagnation, thereby enabling continued improvement during prolonged RL. Consequently, we do not view the inclusion of KL regularization as a limitation of QRPO; rather, we recommend adopting a dynamic-reference approach, where the reference policy is periodically updated to facilitate further reward optimization. Moreover, this strategy helps prevent saturation of the training signal when the trained policy approaches the top quantile, ensuring the quantile reward remains informative throughout training.
@@ Appendix B Guidelines for using QRPO with quantile transformation functions
$t\rule$[-8.0pt]{0.0pt}{24.0pt} $\beta\bigl(e^{1/\beta}-1\bigr)\rule$[-8.0pt]{0.0pt}{24.0pt}
$\log t\rule$[-8.0pt]{0.0pt}{24.0pt} $\dfrac{\beta}{\beta+1}\rule$[-8.0pt]{0.0pt}{24.0pt}
$t^{2}\rule$[-8.0pt]{0.0pt}{24.0pt} $\dfrac{\sqrt{\pi\,\beta}}{2}\,\operatorname{erfi}\bigl(1/\sqrt{\beta}\bigr)\rule$[-8.0pt]{0.0pt}{24.0pt}
$\sqrt{t}\rule$[-8.0pt]{0.0pt}{24.0pt} $2\beta\left[\,\beta+(1-\beta)\,e^{1/\beta}\right]\rule$[-8.0pt]{0.0pt}{24.0pt}
$\texttt{CDF}^{-1}_{\mathcal{N}(0,1)}(t)\rule$[-8.0pt]{0.0pt}{24.0pt} $\exp\bigl(\tfrac{1}{2\beta^{2}}\bigr)\rule$[-8.0pt]{0.0pt}{24.0pt}
$\mu+\sigma\,\texttt{CDF}^{-1}_{\mathcal{N}(0,1)}(t)\rule$[-8.0pt]{0.0pt}{24.0pt} $\exp\bigl(\tfrac{\mu}{\beta}+\tfrac{\sigma^{2}}{2\beta^{2}}\bigr)\rule$[-8.0pt]{0.0pt}{24.0pt}
@@
$\mathbf{\tilde{\mathcal{R}}}(x,y)=f\left(\mathbf{\mathcal{R}}_{q}(x,y)\right).$
@@
The optimal solution for the KL-regularized RL fine-tuning objective (Eq. 2) depends on the shape of the reward distribution; therefore, the primary function of a quantile transformation $f$ is to modify the transformed reward distribution to have some useful properties. One can easily obtain any desired distribution by using its inverse CDF as a quantile transformation function (since for any distribution $Y$, applying its inverse CDF to a uniform sample, $x\sim\texttt{Uniform}[0,1]$, results in $\texttt{CDF}_{Y}^{-1}(x)\sim Y$). For example, one might wish the reward distribution to be Gaussian with mean $\mu$ and variance $\sigma^{2}$. However, after transforming it into a quantile reward $\mathcal{R}_{q}$, the resulting reward is uniformly distributed. In this case, the following transformation can be applied to recover the desired distribution: $\tilde{\mathcal{R}}=f(\mathcal{R}_{q})=\sigma\cdot\texttt{CDF}^{-1}_{\mathcal{N}(0,1)}(\mathcal{R}_{q})+\mu$.
@@
One possible approach to address the finiteness condition is to get insights from the distribution of transformed rewards $\mathbf{\tilde{\mathcal{R}}}$. The partition function $\tilde{Z}(x)$ is simply a moment generating function (MGF) for the distribution of transformed rewards under the reference model: $\tilde{Z}(x)=\mathbb{E}_{\tilde{r}}\left[e^{t\tilde{r}}\right]=M_{\tilde{r}}(t)$, where $t=\frac{1}{\beta}$, $\tilde{r}=\mathbf{\tilde{\mathcal{R}}}(x,y)$ for $y\sim\pi_{ref}(\cdot\mid x)$ (see Section D for the derivation). Thus, to check the finiteness of $\tilde{Z}(x)$, one can derive the distribution of the transformed rewards for completions sampled from the reference model, and then check the value of the corresponding MGF for $t=\frac{1}{\beta}$.
@@
There is a sufficient condition for the finiteness of $\tilde{Z}$: if the transformation function $f$ is upper-bounded on the interval $[0,1]$, then the corresponding partition function is guaranteed to be finite (proof in Appendix G). However, it is very important to note that this condition is not necessary, i.e., there exist some unbounded functions (we consider the upper bound) such that the corresponding $\tilde{Z}$ is finite. An important example of such an unbounded function is the inverse CDF for the normal distribution (the transformation we considered above to obtain a Gaussian distribution of transformed rewards).
@@
After applying a quantile transformation to the initial reward values, the distribution of the resulting quantile reward $\mathcal{R}_{q}$ is known; therefore, the corresponding partition function $Z_{q}$ can be calculated analytically. Moreover, starting from a quantile reward $\mathcal{R}_{q}$, one can apply any additional transformation $f(\mathcal{R}_{q})$ to it while still preserving the ability to compute the corresponding partition function: the partition function corresponding to a specific transformation function $f$ can be computed as a simple integral (derivation in Appendix F):
$\tilde{Z}(x)=\int_{0}^{1}\exp{\left(\frac{1}{\beta}f\left(t\right)\right)}dt.$
@@
We can make a more rigorous observation regarding the correct scaling of $\beta$ in the case where the quantile transformation function $f$ is upper-bounded on the interval $[0,1]$ (see Appendix K for the intuition). The key factor determining the relative scale of the transformed reward is the left derivative of the quantile transformation function $f$ at $x=1$. Therefore, when comparing two upper-bounded quantile transformation functions, the ratio between their respective $\beta$ values should be approximately equal to the ratio of their left derivatives at $x=1$. If both left first derivatives at $x=1$ are equal to zero, move to higher orders until you find the first non-zero left derivative common to both functions. Let $k$ be that order. Then the appropriate $\beta$-scaling factor is the ratio of their $k$-th left derivatives. If the first non-zero left derivative appears at different orders for the two functions, no single constant can scale $\beta$ consistently.
Appendix C Statistical Consistency and Convergence Guarantees
In this section, we analyze the QRPO objective in Eq. (11) (and its practical variant Eq. (12)) through the lens of (i) exactness of the partition function, (ii) statistical properties of the empirical quantile reward, and (iii) optimization with stochastic gradients.
Notation.
Write the calibrated target as $s(x,y)\coloneqq\mathcal{R}_{q}(x,y)-\beta\log Z_{q}$ and the model log–ratio as $g_{\theta}(x,y)\coloneqq\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}.$ The population QRPO loss is $\mathcal{L}(\theta)=\mathbb{E}\big[(s(x,y)-g_{\theta}(x,y))^{2}\big].$
C.1 Exactness of the partition function
Assume that, for each fixed $x$, the reference reward distribution $r=\mathcal{R}(x,y)$ for $y\sim\pi_{\mathrm{ref}}(\cdot\mid x)$ is continuous (as with standard regression reward models). Then the quantile reward $\mathcal{R}_{q}(x,y)=F_{\mathrm{ref}}(x,\mathcal{R}(x,y))$ is uniformly distributed on $[0,1]$ for every $x$. Therefore the partition function is analytically exact and $x$-independent:
$Z_{q}(x)=\mathbb{E}\left[\exp\Big(\tfrac{1}{\beta}\mathcal{R}_{q}\Big)\right]=\int_{0}^{1}e^{t/\beta}\,dt=\beta\left(e^{1/\beta}-1\right).$ (17)
Hence QRPO introduces no approximation or modeling bias through $Z_{q}$.
Remark (discrete rewards)
When $\mathcal{R}$ takes a small number of values, $\mathcal{R}_{q}$ is no longer uniform; however $Z_{q}(x)=\sum_{k}p_{k}(x)\,e^{t_{k}/\beta}$ remains exactly computable as a finite expectation over the atoms $\{t_{k}\}$ with masses $\{p_{k}(x)\}$ (see App. F). Thus $Z_{q}$ is still exact.
Practical constant used in Eq. (12)
If one replaces $\beta\log Z_{q}$ by $\beta\log\beta+1$ for stability, the calibration error is $\Delta_{\beta}=\beta\log Z_{q}-(\beta\log\beta+1)=\beta\log(1-e^{-1/\beta})<0.$ For $\beta\leq 0.1$, $|\Delta_{\beta}|\leq 2\beta e^{-1/\beta}<10^{-5}$, is negligible.
C.2 Empirical quantile reward: unbiasedness and concentration
For each $x$, let $y_{1},\dots,y_{n}\sim\pi_{\mathrm{ref}}(\cdot\mid x)$ be i.i.d. reference samples with rewards $r_{i}=\mathcal{R}(x,y_{i})$, and define the empirical quantile estimator $\widehat{\mathcal{R}}_{q}(x,y)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{r_{i}\leq\mathcal{R}(x,y)\}.$
[…]
Uniform consistency and finite-sample deviation By Glivenko–Cantelli, $\sup_{r}|\widehat{F}_{n}(x,r)-F_{\mathrm{ref}}(x,r)|\xrightarrow[]{\text{a.s.}}0$ as $n\to\infty$. Moreover, the Dvoretzky–Kiefer–Wolfowitz inequality yields, for any $\varepsilon>0$, the following distribution-free bound:
$\Pr\left(\,|\widehat{\mathcal{R}}_{q}(x,y)-\mathcal{R}_{q}(x,y)|>\varepsilon\ \middle|\ x\right)\ \leq\ 2\,e^{-2n\varepsilon^{2}}.$
Expected error
$\mathrm{Var}\left(\widehat{\mathcal{R}}_{q}(x,y)\mid x,y\right)=\frac{\mathcal{R}_{q}(x,y)\bigl(1-\mathcal{R}_{q}(x,y)\bigr)}{n}\leq\frac{1}{4n},\qquad\Rightarrow\ \mathbb{E}\left|\widehat{\mathcal{R}}_{q}-\mathcal{R}_{q}\right|\leq\frac{1}{2\sqrt{n}}.$
Thus the only stochasticity introduced by QRPO is zero-mean label noise on the target, with sub-Gaussian tails (from the exponential bound) and $O(1/\sqrt{n})$ rate.
C.3 Population argmin is unchanged by unbiased quantile noise
Let $\xi(x,y)\coloneqq\widehat{\mathcal{R}}_{q}(x,y)-\mathcal{R}_{q}(x,y)$, so that $\mathbb{E}[\xi(x,y)\mid x,y]=0$. Define $\widehat{s}(x,y)=s(x,y)+\xi(x,y)$, where $s(x,y)=\mathcal{R}_{q}(x,y)-\beta\log Z_{q}$ and $g_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}$. Then the population loss when training against $\widehat{s}$ is
$\widehat{\mathcal{L}}(\theta)=\mathbb{E}\left[\big(\widehat{s}(x,y)-g_{\theta}(x,y)\big)^{2}\right]=\mathcal{L}(\theta)+\mathbb{E}\left[\xi(x,y)^{2}\right],$
because the cross term vanishes: $\mathbb{E}\left[\xi(x,y)\,\big(s(x,y)-g_{\theta}(x,y)\big)\right]=\mathbb{E}\left[\mathbb{E}\left[\xi(x,y)\mid x,y\right]\big(s-g_{\theta}\big)\right]=0.$
Oracle risk invariance The set of minimizers of $\widehat{\mathcal{L}}(\theta)$ equals that of $\mathcal{L}(\theta)$. Unbiased quantile noise adds only a $\theta$–independent constant to the population risk.
Gradient perturbation bound (per sample)
With the same $\xi(x,y)$ as above, the per-sample gradient difference is
$\nabla_{\theta}\widehat{\ell}-\nabla_{\theta}\ell=2\,\xi(x,y)\,\nabla_{\theta}g_{\theta}(x,y).$
If $\|\nabla_{\theta}g_{\theta}(x,y)\|\leq G$ (e.g., via gradient clipping), then for any $\varepsilon>0$,
$\Pr\left(\|\nabla_{\theta}\widehat{\ell}-\nabla_{\theta}\ell\|\leq 2G\,\varepsilon\ \middle|\ x\right)\ \geq\ 1-2e^{-2n\varepsilon^{2}}.$
Equivalently, with probability at least $1-\delta$, $\|\nabla_{\theta}\widehat{\ell}-\nabla_{\theta}\ell\|\leq 2G\,\sqrt{\tfrac{1}{2n}\log\tfrac{2}{\delta}}.$
C.4 Convergence of (stochastic) gradient descent
The QRPO loss is nonconvex in $\theta$, but the preceding results ensure that the population objective is unaffected by empirical-quantile noise. Assume (A1) the loss is lower-bounded and $L$-smooth in $\theta$; (A2) stochastic gradients have bounded second moment; (A3) either the empirical quantiles are resampled independently across iterations (or $n\to\infty$ so that $\xi\to 0$), or one optimizes the fixed empirical objective with standard SGD. Then:
  1. 1.
    With resampling (or $n\to\infty$), SGD produces gradients that are unbiased for $\nabla\mathcal{L}(\theta)$ and converges to a stationary point of the oracle loss $\mathcal{L}(\theta)$ under standard step-size conditions $\sum_{t}\eta_{t}=\infty,\ \sum_{t}\eta_{t}^{2}<\infty$.
  2. 2.
    With fixed precomputed $\widehat{\mathcal{R}}_{q}$ (Algorithm 5), SGD converges to a stationary point of the empirical QRPO loss. By DKW and Glivenko–Cantelli, as $n\to\infty$ the empirical objective converges uniformly to the oracle objective, so the stationary points approach those of $\mathcal{L}(\theta)$.
Summary.
(i) $Z_{q}$ is known in closed form (or as an exact finite sum in the discrete case), so QRPO introduces no bias via the partition function; the practical constant in Eq. (12) differs from $\beta\log Z_{q}$ by a negligible $O(\beta e^{-1/\beta})$ shift. (ii) The empirical quantile reward is an unbiased estimator with exponential concentration $e^{-2n\varepsilon^{2}}$ and expected error $O(1/\sqrt{n})$. (iii) This induces only zero-mean label noise, leaving the population argmin unchanged; SGD converges to a stationary point of the same oracle objective under standard assumptions.
Appendix D The partition function is a moment generating function at $t=1/{\beta}$